OpenStudy (anonymous):

the product of 3 and a number increased by 8 is -48 what is the number?

OpenStudy (anonymous):

"the product of 3 and a number" indicated multiplication and "increased by 8 " indicates addition, and "is -48" indicates what it all equals. So, we can set up an equation like this: 3x+8= -48 Do you know how to solve for x?

OpenStudy (anonymous):

is it 18 and 2/3?

OpenStudy (anonymous):

@Butterfly16

OpenStudy (anonymous):

Well, lets look: 3x+8= -48 First, we're going to subtract 8 from both sides giving us: 3x= -56 Now, we're going to divide both sides by 3 to get that x alone: 3x= -56 /3 /3 x= -18.6

OpenStudy (anonymous):

alright thanks, was just checking

OpenStudy (anonymous):

Anytime :o)